Newman Spear invested $18,300 today in a fund that earns 12% compounded annually. Click here to view factor tables To what amount will the investment grow in 4 years? To what amount would the investment grow in 4 years if the fund earns 12% annual interest compounded semiannually? (Round factor values to 5 decimal places, e.g. 1.25124 and final answers to 0 decimal places, e.g. 458,581.) Investment at 12% annual interest $enter a dollar amount rounded to 0 decimal places Investment at 12% annual interest, compounded semiannually $enter a dollar amount rounded to 0 decimal places
Answer #1

1.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$18300*(1.12)^4

=18300*1.57352

=$28795(Approx).

2.We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$18300*(1+0.12/2)^(2*4)

$18300*1.59385

=$29167(Approx).